Question 2 (2 points) Which of the following statement/ statements is not an example of agency problem? i. The managers buy expensive dinners on company account too often i. The company introduces stock option plans for managers ili. The company spends a hefty amount on market research i and ii only i and ili only i and ii only i. ii and ii Question 3 (2 points) What is the PV of $5325 to be received in one year from today if the interest rate is 6.50% compounded annually? $5,000 $5,07143 $5,671.13 $5,591.25
Answer #1

(2) Agency Problem in a relationship is a conflict of interest which happens when one party is expected to act in the best interest of the other party.

Hence, in (ii) and (iii), there is no conflict of interest since these activities are done either to reward the employees or to improve companies revenue.

Hence, option (b) ii and iii only is the correct option

(3) Given Future Value (FV) = $5325

n = 1 year

r = interest rate = 6.50%

Hence, PV = FV / (1+r)n = 5325 / (1.065) = $5000

Option (a) $5000 is the correct answer